OpenStudy (anonymous):

Enter the variable that is easiest to solve for in this system of equations. y=5x+5 3y=5x+7

OpenStudy (anonymous):

x: Since you already have y in terms of x simply plug in y=5x+5 into the second equation. 3(5x+5)=5x+7 then solve for x.

OpenStudy (anonymous):

y. you can use elimination and simply get rid of all the "x"s and be left with an equation of y -y=-5x-5 3y=5x+7 when we add these vertically: 2y=2 y=1 then we can solve for x